# Regular Kerr black holes: The fellowship of the string

### v1

@@ III The matching at the ring
Even though the Darmois-Israel formalism (DIF) [27] for matching different metrics is well known, we review the main aspects of such a formalism here in order to set up notation and to present the important quantities adapted to the present problem.
@@
Let us then refer the interior disk $\mathcal{D}$ with the metric given by Eq. (8) as the spacetime $(\mathcal{M}_{i},\,g_{i})$, and to the equatorial region outside the ring $\mathcal{A}$ with the metric given by Eq. (13) in a first study, and by (11) in a second study, as the spacetime $(\mathcal{M}_{e},\,g_{e})$. These two spacetimes are to be matched together at the ring $\mathcal{S}$. Let $\xi^{a}=(t,\varphi)$ be the intrinsic coordinates of the ring $\mathcal{S}$, whose metric is given by Eq. (9). Without lost of generality, we already identified the coordinates $t$ and $\varphi$ in the three regions $\mathcal{M}_{i}$, $\mathcal{M}_{e}$, and $\mathcal{S}$.
@@
The DIF deals with the first and the second fundamental forms on $\mathcal{S}$, $h_{ab}$ and $K_{ab}$, respectively. Such quantities may be defined in terms of the geometric quantities in the spacetimes $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ by relations of the form $h_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}g_{\mu\nu}$ and $K_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}\nabla_{\nu}n_{\mu}$, where $e_{a}^{\ \mu}=\partial x^{\mu}/\partial\xi^{a}$, $n_{\mu}$ is the spacelike unit normal vector to the ring $\mathcal{S}$, and $\nabla_{\nu}$ stands for the covariant derivative compatible with the Lorentzian metric. There is one of such relations for each one of the spacetimes $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$.
@@
Moreover, in the DIF, it is necessary to choose the orientation of the unit normal vector to the boundary between the two spacetime regions, $n_{\mu}$, which we choose pointing from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$. The normal to a closed thin shell may have two distinct orientations relative to the center of the coordinates of each spacetime. In other words, the orientation of the normal with respect to the respective radial coordinate of each spacetime is not fixed a priory. Understanding this fact is essential for uncovering all possible thin shell solutions for the present problem.
@@
The first boundary condition in the DIF is the continuity of the first fundamental form across the boundary $\mathcal{S}$, i.e., $[h_{ab}]=0$. Here we employ the standard notation $[Q]\equiv Q\big{|}^{e}_{\mathcal{S}}-Q\big{|}^{i}_{\mathcal{S}}$, with $Q\big{|}^{e}_{\mathcal{S}}$ denoting any given quantity $Q$ evaluated on $\mathcal{S}$ from the perspective of region $\mathcal{M}_{e}$, and similarly for $Q\big{|}^{i}_{\mathcal{S}}$. It is easy to see that, in the present case, the induced metric $h_{ab}$ at $\mathcal{S}$ as perceived from each spacetime is the same.
@@
The second boundary condition establishes the relation between the possible discontinuity of the second fundamental form $K_{ab}$ across the boundary $\mathcal{S}$ and the energy-momentum (EMT) of the boundary layer. I.e., if $[K_{ab}]\neq 0$ then a thin shell at $\mathcal{S}$ is needed, and the respective energy-momentum tensor $S_{ab}$ is given by
@@ IV.1 The interior region
$n_{i\mu}=\epsilon_{i}\delta_{\mu}^{\rho},$ (15)
where $\epsilon_{i}$ is a normalization factor. The normalization condition $g^{\mu\nu}_{i}n_{i\mu}n_{i\nu}=1$ with metric (9) implies in $\epsilon_{i}=\pm 1$. The sign $\epsilon_{i}$ is the quantity that determines the relative orientation of $n_{i\mu}$ in relation to the coordinate $\rho$, as one can see from the scalar product between $n_{i\mu}$ and the gradient of the $\rho$, i.e., $\nabla_{\mu}\rho$. For $\epsilon_{i}=+1$, we have $g^{\mu\nu}_{i}n_{i\mu}(\nabla_{\nu}\rho)=+1>0$, which implies that the normal points in the direction of increasing $\rho$ as seen from $\mathcal{M}_{i}$. On the other hand, for $\epsilon_{i}=-1$, we have $g^{\mu\nu}_{i}n_{i\mu}(\nabla_{\nu}\rho)=-1<0$, which implies that the normal points in the direction of decreasing $\rho$ as seen from $\mathcal{M}_{i}$.
For $\rho=a$ the metric (8) defined in $\mathcal{M}_{i}$ reduces to the metric (9) defined in $\mathcal{S}$, and then the extrinsic curvature of $\mathcal{S}$ as seen from $\mathcal{M}_{i}$ is given by
$K^{t}_{it}=0,\ \ K^{t}_{i\varphi}=0=\ K^{\varphi}_{it},\ \ K^{\varphi}_{i\varphi}=\frac{\epsilon_{i}}{a}.$ (16)
@@ IV.2 The exterior region
$n_{\mu}^{+}=\epsilon_{e}\delta_{\mu}^{\varrho},$ (17)
where $\epsilon_{e}$ is the normalization factor. The condition $g^{\mu\nu}_{e}n_{e\mu}n_{e\nu}=1$ with metric (13) implies that $\epsilon_{e}=\pm 1$ and, therefore, determines the sign of $n_{e\mu}$. The sign $\epsilon_{e}$ is the quantity that determines the relative orientation between $n_{e\mu}$ and the coordinate $\varrho$. For $\epsilon_{e}=+1$, we have $g^{\mu\nu}_{e}n_{e\mu}(\nabla_{\nu}\varrho)=+1>0$, which implies that the normal points in the direction of increasing $\varrho$ as seen from $\mathcal{M}_{e}$. Whereas for $\epsilon_{i}=-1$, we have $g^{\mu\nu}_{e}n_{e\mu}(\nabla_{\nu}\varrho)=-1<0$, which implies that the normal points in the direction of decreasing $\varrho$ as seen from $\mathcal{M}_{e}$. Since, for $\varrho=a$, the metric (13) of $\mathcal{M}_{e}$ reduces to the metric (9) of $\mathcal{S}$, then the extrinsic curvature of $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is given by
$K^{t}_{et}=-2m_{0}a\epsilon_{e},\ \ K^{t}_{e\varphi}=0=\ K^{\varphi}_{et},\ \ K^{\varphi}_{e\varphi}=\frac{\epsilon_{e}}{a}.$ (18)
@@ IV.3 Are all thin shell solutions possible?
IV.3.1 The matter-energy content at the ring
The Darmois-Israel formalism allows for both values of $\epsilon_{i}$ and $\epsilon_{e}$, which lead, in principle, to four possible thin shell solutions. However, the particular choices of $\epsilon_{i}$ and $\epsilon_{e}$ determine the possible domains of the coordinates $\rho$ and $\varrho$ which, at the end, have to be compared with the domains previously established for such coordinates.
@@ IV.3.1 The matter-energy content at the ring
are the energy-density and the pressure of the fluid, respectively, $h_{ab}$ is the metric of the ring given in Eq. (9), and $u_{a}=\delta_{a}^{\ t}$ is the fluid $2$-velocity at the ring. Hence, differently from the interpretation given in Ref. [7], we can see from the form of the fluid’s velocity that the string is not rotating.
Let us now look at the four possible different cases.
@@ IV.3.2 The case $\epsilon_{i}=+1$ and $\epsilon_{e}=-1$
By starting with the choice made by Smailagic and Spallucci [7], i.e., $\epsilon_{i}=+1$ and $\epsilon_{e}=-1$, we get the linear energy density and the pressure of the string as $\sigma=1/4\pi a$ and $p=m_{0}a/4\pi$, respectively. The linear energy density is always positive, while the pressure can be positive or negative depending on the sign of $m_{0}$.
More importantly, let us notice that $\epsilon_{i}=+1$ implies that the normal vector points in the direction of increasing radial coordinate as seen from the interior region. Since the coordinate $\rho$ is in the interval $\rho\in[0,\,a)$, and recalling our convention that the normal vector $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the choice $\epsilon_{i}=+1$ implies that there is a center in the interior region, as expected. In turn, the choice $\epsilon_{e}=-1$ implies that the normal vector points in the direction of the decreasing radial coordinate as seen from the exterior region. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, this implies that in the exterior region one also has $\varrho\in[0,\,a]$. Hence, this possible solution has two centers, one in each region, which implies that we have a kind of Minkowski-dS or Minkowski-AdS closed universe that can be interpreted as a bubble universe, as similarly done in Ref. [29] for a Minkowski-Minkowski closed universe. However, the original domain of the coordinate $\varrho$ is assumed to be $\varrho\in[a,a+\varepsilon)$, for $\varepsilon>0$, with the already established notion that the exterior region corresponds to the equatorial plane just outside the ring. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
@@ IV.3.3 The case $\epsilon_{i}=-1$ and $\epsilon_{e}=+1$
The other non trivial case is given by the choice $\epsilon_{i}=-1$ and $\epsilon_{e}=+1$. In this case, the linear energy density and the pressure of the fluid are given by $\sigma=-1/4\pi a$ and $p=-m_{0}a/4\pi$, respectively. The linear energy density is negative, while the pressure can be positive or negative depending on the sign of $m_{0}$.
More importantly, let us notice that $\epsilon_{i}=-1$ implies that the normal points in the direction of decreasing radial coordinate as seen from the interior region, while $\epsilon_{e}=+1$ implies that the normal points in the direction of increasing radial coordinate as seen from the exterior region. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, this implies that in both regions the radial coordinates start from $a$ and go to infinity. Therefore, the possible solution has no center at all and corresponds to a kind of Minkowski-dS or Minkowski-AdS open universe and can be interpreted as a traversable wormhole, as similarly done in Ref. [29] for a Minkowski-Minkowski open universe. However, once more let us notice that the original domain of the interior radial coordinate is $\rho\in[0,a]$, with the already established notion of the interior region corresponding to the region inside the ring, i.e., corresponding to the disk. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
@@ IV.3.4 The case $\epsilon_{i}=-1$ and $\epsilon_{e}=-1$
For the choice $\epsilon_{i}=\epsilon_{e}=-1$, the linear energy density and the pressure of the fluid are given by $\sigma=0$ and $p=m_{0}a/4\pi$, respectively. The linear energy density is identically zero, while the pressure can be positive or negative depending on the sign of $m_{0}$.
In this case, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ are basically swapped. More importantly, let us notice that this choice $\epsilon_{i}$ and $\epsilon_{e}$ implies that the normal points in the direction of decreasing radial coordinate as seen from both regions. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the role of the radial coordinates $\rho$ and $\varrho$ are swapped. However, this contradicts the original domains of the radial coordinates and the notions that the interior region corresponds to the interior disk and the exterior region corresponds to the equatorial disk just outside the ring. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.

### v2

@@ III The matching at the ring
Although the Darmois-Israel formalism (DIF) [31] for matching different metrics is well known, we review the main aspects of such a formalism here in order to set up notation and present the important quantities adapted to the present problem.
@@
Let us then refer to the interior disk $\mathcal{D}$ with the metric given by Eq. (19) as the spacetime $(\mathcal{M}_{i},\,g_{i})$, and to the equatorial region outside the ring $\mathcal{A}$ with the metric given by Eq. (25) in a first study and by (24) in a second study, as the spacetime $(\mathcal{M}_{e},\,g_{e})$. These two spacetimes are to be matched at the ring $\mathcal{S}$. Let $\xi^{a}=(t,\varphi)$ be the intrinsic coordinates of the ring $\mathcal{S}$, whose metric is given by Eq. (20). Without loss of generality, we have already identified the timelike coordinate $t$ and the azimuthal coordinate $\varphi$ in the three regions $\mathcal{M}_{i}$, $\mathcal{M}_{e}$, and $\mathcal{S}$.
@@
The DIF deals with the first and second fundamental forms on $\mathcal{S}$, $h_{ab}$ and $K_{ab}$, respectively. Such quantities may be defined in terms of geometric quantities in the spacetimes $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ by relations of the form $h_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}g_{\mu\nu}$ and $K_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}\nabla_{\nu}n_{\mu}$, where $e_{a}^{\ \mu}=\partial x^{\mu}/\partial\xi^{a}$, $n_{\mu}$ is the spacelike unit vector normal to the ring $\mathcal{S}$, and $\nabla_{\nu}$ stands for the covariant derivative compatible with the Lorentzian metric. There is one such relation with a corresponding quantity for each one of the spacetime regions $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$. For example, the unit normal vector $n_{\mu}$ has two representations: $n_{(i)u}$ is the unit normal vector as seen from the point of view of the internal section $\mathcal{M}_{i}$, while $n_{(e)\mu}$ is the unit normal vector as seen from the point of view of the external section $\mathcal{M}_{e}$.
@@
Moreover, in the DIF, it is necessary to choose the orientation of the unit normal vector to the boundary between the two spacetime regions ($n_{(i)\mu}$ and $n_{(e)\mu}$), which we choose to point from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$. The normal to a closed thin shell may have two distinct behaviors relative to the center of the coordinates in each spacetime, i.e., its orientation regarding the increasing or decreasing of the radial coordinate of each spacetime is not fixed a priori [34]. This implies that different sections of the interior and exterior spacetimes bounded by the ring can be glued to each other. Understanding this fact is essential for uncovering all possible thin shell solutions for the present problem.
@@
The first boundary condition in the DIF is the continuity of the first fundamental form across the boundary $\mathcal{S}$, i.e., $[h_{ab}]=0$. Here we employ the standard notation $[Q]\equiv Q\big{|}^{e}_{\mathcal{S}}-Q\big{|}^{i}_{\mathcal{S}}$, with $Q\big{|}^{e}_{\mathcal{S}}$ denoting any given quantity $Q$ evaluated on $\mathcal{S}$ from the perspective of region $\mathcal{M}_{e}$, and similarly for $Q\big{|}^{i}_{\mathcal{S}}$. It is easy to see that, in the present case, the induced metric $h_{ab}$ on $\mathcal{S}$ as perceived from each spacetime region is the same and, with our choice of coordinates, is given by Eq. (20).
@@
The second boundary condition establishes the relation between the possible discontinuity of the second fundamental form $K_{ab}$ across the boundary $\mathcal{S}$ and the energy-momentum tensor (EMT) of the boundary layer. I.e., if $[K_{ab}]\neq 0$ then a thin shell at $\mathcal{S}$ is needed, and the respective energy-momentum tensor $S_{ab}$ is given by
@@ IV.1 The interior region
$n_{(i)\mu}=\epsilon_{i}\delta_{\mu}^{\rho},$ (27)
where $\epsilon_{i}$ is a normalization factor. The normalization condition $g^{\mu\nu}_{i}n_{(i)\mu}n_{(i)\nu}=1$ with the metric (20) implies $\epsilon_{i}=\pm 1$. The sign $\epsilon_{i}$ is the quantity that determines the relative orientation of $n_{(i)\mu}$ in relation to the coordinate $\rho$, as one can see from the scalar product between $n_{(i)\mu}$ and the gradient of $\rho$, i.e., $\nabla_{\mu}\rho$. For $\epsilon_{i}=+1$, we have $g^{\mu\nu}_{i}n_{(i)\mu}(\nabla_{\nu}\rho)=+1>0$, which implies that the normal points in the direction of increasing $\rho$ as seen from $\mathcal{M}_{i}$. On the other hand, for $\epsilon_{i}=-1$, we have $g^{\mu\nu}_{i}n_{(i)\mu}(\nabla_{\nu}\rho)=-1<0$, which implies that the normal points in the direction of decreasing $\rho$ as seen from the point of view of an observer in $\mathcal{M}_{i}$.
[…]
$K^{t}_{(i)t}=0,\ \ K^{t}_{(i)\varphi}=0=\ K^{\varphi}_{(i)t},\ \ K^{\varphi}_{(i)\varphi}=\frac{\epsilon_{i}}{a}.$ (28)
Obviously, the frame-dragging velocity in the interior region vanishes since the $g_{\varphi t}$ component of the interior metric (19) is identically zero.
@@ IV.2 The exterior region
$n_{(e)\mu}=\epsilon_{e}\delta_{\mu}^{\varrho},$ (29)
[…]
$\!\!K^{t}_{(e)t}=-2m_{0}a\epsilon_{e},\;K^{t}_{(e)\varphi}=0=\ K^{\varphi}_{(e)t},\;K^{\varphi}_{(e)\varphi}=\frac{\epsilon_{e}}{a}.$ (30)
As a final comment at this point, we observe that the frame-dragging velocity in the exterior region close to the string vanishes, since the $g_{\varphi t}$ metric component of the exterior metric (25) is identically zero.
@@ IV.3 Are all thin shell solutions possible?
IV.3.1 The matter-energy content of the ring
The Darmois-Israel formalism allows for both values of $\epsilon_{i}$ and $\epsilon_{e}$, which, in principle, leads to four possible thin shell solutions. However, the particular choices of $\epsilon_{i}$ and $\epsilon_{e}$ determine the possible domains of the coordinates $\rho$ and $\varrho$ that, in the end, must be in accordance with the original domain previously established for the Boyer-Lindquist radial coordinate, $0\leq r<\infty$.
@@ IV.3.1 The matter-energy content of the ring
are the energy density and the pressure (or tension) of the fluid of the ring, respectively, $h_{ab}$ is the metric of the ring given in Eq. (20), and
$u_{a}=\delta_{a}^{\ t},$ (33)
is the fluid velocity at the ring.
Notice that the velocity (33) does not carry a component along the azimuthal direction $\varphi$, which implies that the string fluid does not rotate with respect to the stationary observer at infinity. In Ref. [8], the equatorial angular velocity of the Gürses-Gürsey fluid in the limit $r=0$, $\Omega=1/a$ [see Eq. (12)], is attributed to the thin shell itself. However, this assumption is not appropriate once the Gürses-Gürsey background and the thin shell have different matter contents, as it can be seen from the different energy-momentum tensors specified by Eqs. (8) and (32), respectively.
In addition to the vanishing of the fluid angular velocity, as we comment in Sec. II.1, the frame-dragging of the non-singular Kerr metric vanishes at the ring $\cal S$ as well, a result that is also consistent with the approximate metric (25), so that no rotational effect is observed at the string.
Let us now look at the four possible different cases for the choices of $\epsilon_{i}$ and $\epsilon_{e}$.
@@ IV.3.2 The case $\epsilon_{i}=+1$ and $\epsilon_{e}=-1$
Starting with the choice made by Smailagic and Spallucci [8], i.e., $\epsilon_{i}=+1$ and $\epsilon_{e}=-1$, we get the line energy density and the pressure of the string as $\sigma=1/4\pi a$ and $p=m_{0}a/4\pi$, respectively. The line energy density is always positive, while the pressure can be positive or negative depending on the sign of $m_{0}$.
Figure 1: Embedding diagram of a $t=\text{constant}$ and $\theta=\pi/2$ slice of $(\mathcal{M}_{i},g_{i})$ with $\epsilon_{i}=+1$ and $(\mathcal{M}_{e},g_{e})$ with $\epsilon_{e}=-1$ in $3$-dimensional Euclidean space. The domain of the interior radius is $0\leq\rho\leq a$. The domain of the exterior radius is also $0\leq\varrho\leq a$. The string is represented by the black solid lines in each spacetime, with the lines being identified through the junction.
More importantly, let us notice that $\epsilon_{i}=+1$ implies that the normal vector points in the direction of increasing radial coordinate, as seen from the interior region. Since the coordinate $\rho$ is in the interval $\rho\in[0,\,a)$, and recalling our convention that the normal vector $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the choice $\epsilon_{i}=+1$ implies that there is a center in the interior region, as expected. In turn, the choice $\epsilon_{e}=-1$ implies that the normal vector points in the direction of decreasing radial coordinate, as seen from the exterior region. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, this implies that the exterior region is also a disk with radius $a$, $\varrho\in[0,\,a]$. Hence, as it can be seen from Fig. 1, this possible solution has two centers, one in each region, which means that a disk of the Minkowski (interior) geometry is glued to another disk-shaped section of the exterior geometry. The resulting space is a kind of Minkowski-dS or Minkowski-AdS closed universe that can be interpreted as a bubble universe, as is similarly done in Ref. [34] for a Minkowski-Minkowski closed universe. However, the original domain of the coordinate $\varrho$ is $\varrho\in[a,a+\varepsilon)$, for $\varepsilon>0$, with the already established notion that the exterior region corresponds to the equatorial plane just outside the ring. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
@@ IV.3.3 The case $\epsilon_{i}=-1$ and $\epsilon_{e}=+1$
Another nontrivial case is given by the choice $\epsilon_{i}=-1$ and $\epsilon_{e}=+1$. In this case, the line energy density and the pressure of the fluid are given by $\sigma=-1/4\pi a$ and $p=-m_{0}a/4\pi$, respectively. The line energy density is negative, while the pressure can be positive or negative depending on the sign of $m_{0}$.
Figure 2: Embedding diagram of a $t=\text{constant}$ and $\theta=\pi/2$ slice of $(\mathcal{M}_{i},g_{i})$ with $\epsilon_{i}=-1$ and $(\mathcal{M}_{e},g_{e})$ with $\epsilon_{e}=+1$ in $3$-dimensional Euclidean space. The domain of the interior radius is $a\leq\rho<\infty$. The domain of the exterior radius is also $a\leq\varrho<\infty$. The string is represented by black solid lines in each spacetime, with the lines being identified through the junction.
More importantly, the choice $\epsilon_{i}=-1$ implies that the normal points in the direction of decreasing radial coordinate as seen from the interior region, while $\epsilon_{e}=+1$ implies that the normal points in the direction of increasing radial coordinate as seen from the exterior region. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, this implies that in both regions the radial coordinates start from $a$ and go to infinity. Therefore, the possible solution has no center at all, as it can be seen from Fig. 2, which means that the asymptotic section of the Minkowski (interior) geometry, from which a disk of radius $r=a$ has been removed, is being glued to the asymptotic section of the exterior geometry. The resulting space is a kind of Minkowski-dS or Minkowski-AdS open universe and can be interpreted as a traversable wormhole, as is similarly done in Ref. [34] for a Minkowski-Minkowski open universe. However, once more let us notice that the original domain of the interior radial coordinate is $\rho\in[0,a]$, with the already established notion of the interior region corresponding to the region inside the ring, i.e., corresponding to the disk. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
@@ IV.3.4 The case $\epsilon_{i}=-1$ and $\epsilon_{e}=-1$
For the choice $\epsilon_{i}=\epsilon_{e}=-1$, the line energy density and the pressure of the fluid are given by $\sigma=0$ and $p=m_{0}a/4\pi$, respectively. The line energy density is identically zero, while the pressure can be positive or negative depending on the sign of $m_{0}$.
Figure 3: Embedding diagram of a $t=\text{constant}$ and $\theta=\pi/2$ slice of $(\mathcal{M}_{i},g_{i})$ with $\epsilon_{i}=-1$ and $(\mathcal{M}_{e},g_{e})$ with $\epsilon_{e}=-1$ in $3$-dimensional Euclidean space. The domain of the interior radius is $a\leq\rho<\infty$. The domain of the exterior radius is $0\leq\varrho\leq a$. The string is represented by the black solid lines in each spacetime, with the lines being identified through the junction.
In this case, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ are basically swapped. More importantly, note that this choice for $\epsilon_{i}$ and $\epsilon_{e}$ implies that the normal points in the direction of decreasing radial coordinates, as seen from both regions. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the role of the radial coordinates $\rho$ and $\varrho$ are swapped. This means that a disk of the de Sitter (exterior) geometry is being glued to the asymptotic section of the Minkowski (interior) geometry from which a disk of radius $r=a$ has been removed, as illustrated in Fig. 3. The resulting space is a kind of dS-Minkowski or AdS-Minkowski universe that presents a central disk that contains a dS(AdS)-type of fluid, and is void of matter everywhere outside the disk. However, this contradicts the original domains of the radial coordinates and the notion that the interior region corresponds to the interior disk and the exterior region corresponds to the equatorial disk just outside the ring. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.